A 1550 kg car is driving north at 37.8 m/s, while another car with a mass of 2681 kg is driving south at 24.5 m/s. What is the total momentum of the system described?
1
Expert's answer
2019-12-23T11:27:26-0500

Momentum of the first car


"p_1=m_1v_1=1550\times 37.8=58590\:\rm kg\cdot m/s\: (north)"

Momentum of the second car


"p_2=m_2v_2=2681\times 24.5=65685\:\rm kg\cdot m/s\; (south)"

The total momentum


"p=p_2-p_1=65685-58590=7095\:\rm kg\cdot m/s\: (south)"